OpenStudy (yamyam70):

x >_ 4 (greater than equal) is same as (infi. , +4 ] , right ?

OpenStudy (anonymous):

The interval should be written the other way around: \(\large [4,\infty)\). Note that \(\large x\geq 4\implies 4\leq x<\infty\). Thus, whenever you write an interval, the smaller (or most negative) value should go first. I hope this clarifies things! :-)

OpenStudy (yamyam70):

Thank you , and good like with your medals ^.^